How do you determine whether an ordered pair is a solution of
an equation?

3x+2y=6 find at least two ordered pairs solutions to this equation.

Answer by Edwin McCravy(20054)

How do you determine whether an ordered pair
is a solution of an equation?
1. Substitute the first coordinate of the ordered pair for x (or the
independent variable) into the equation.

2. Substitute the second coordinate of the ordered pair for y (or the 
independent variable).

3. Simplify each side if necessary so that there is a single number on 
each side of the equation.

4. If the number on the left side is the same number ans the number on
the right side, then the ordered pair is a solution of the equation. If
not, then the ordered pair is not a solution to the equation.   


3x+2y=6 find at least two ordered pairs solutions to this equation.
1. Choose any number arbitrarily. 

I will arbitrarily choose -4 

2. Choose one of the two variables in the equation to substitute it for.

I will arbitrarily choose to substitute -4 for the independent variable x.

3.  Substitute that arbitrarily selected number for that arbitrarily 
selected variable.

   3x+2y=6
3(-4)+2y=6

4. Solve for the other remaining variable.

  -12+2y=6
  +12    +12
--------------
      2y=18
       y=9

5.  Write a left parentheses, "("

    I will write (

6.  Just to the right of the parentheses, write the number you chose 
    in step 1, or that you got in step 4, whichever was the value of
    x, or the independent variable.

    Since I chose -4 for x in step 1, so far I have (-4

7.  Just to the right of that, write a comma, ","

    So far I have (-4,

7.  Just to the right of the comma, write the number you chose 
    in step 1, or that you got in step 4, whichever was the value of
    y, or the dependent variable.  
    
    So far I have (-4,9

8.  Just to the right of the number written in step 7, write a right
    parenthesis, ")". That completes the ordered pair.

    So the ordered pair is (-4,9)

---------------------------

1. Choose any number arbitrarily. 

I will arbitrarily choose 5 

2. Choose one of the two variables in the equation to substitute it for.

I will arbitrarily choose to substitute 6 for the dependent variable y.

3.  Substitute that arbitrarily selected number for that arbitrarily 
selected variable.

  3x+2y=6
3x+2(5)=6


4. Solve for the other remaining variable.

  3x+10=6
    -10 -10
------------
  3x   =-4
   x   =-4/3

5.  Write a left parentheses, "("

    I will write (

6.  Just to the right of the parentheses, write the number you chose 
    in step 1, or that you got in step 4, whichever was the value of
    x, or the independent variable.

    Since I got -4/3 for x in step 4, so far I have (-4/3

7.  Just to the right of that, write a comma, ","

    So far I have (-4/3,

7.  Just to the right of the comma, write the number you chose 
    in step 1, or that you got in step 4, whichever was the value of
    y, or the dependent variable.  
    
    So far I have (-4/3,5

8.  Just to the right of the number written in step 7, write a right
    parenthesis, ")". That completes the ordered pair.

    So the ordered pair is (-4/3,5)

Edwin